OpenStudy (anonymous):

Solve algebraically: y+5=2x^2 -4x 3x-y=1

OpenStudy (anonymous):

put y=3x-1 in the second equation and then solve the quadratic

OpenStudy (anonymous):

2x^2 -4x=3x-1+5 2x^2-7x -4=0 2x^2-8x+x-4=0 2x(x-4) +1(x-4)=0 (2x+1)(x-4)=0 x=-1/2 or x=4

OpenStudy (anonymous):

when x=4 y=3(4)-1=11 when x=-1/2 y=3(-1/2) -1=-5/2

OpenStudy (anonymous):

i set the two equations equal to each other after getting the y's on one side.... I got -1 and -4 as roots..... ?????

OpenStudy (anonymous):

my calculation may not be correct